1. Consider the following simple model of an economy operating with fixed wages, prices, and interest rates, and excess capacity (operating below the full-employment level): where I=900
C=600+0.6Yd+I
Where C is consumption, Yd is disposable income (equal to national income, Y, in absence of a government sector), and I is investment.
a) If aggregate expenditures (AE) are a function of Y, calculate the equilibrium level of income.
b) Calculate the value of the multiplier.
1
Expert's answer
2012-08-14T09:11:12-0400
a) AE =C+I = 600 +0.6Yd +900= 1500 +0.6Yd
b) Multiplier 1/(1-MPC)= 1/0.6=1.67